# Versatile Motion-Language Models for Multi-turn Interactive Agents

### v2

@@ A.13 Prompts for Data collection in Inter-$\text{MT}^{2}$
You are an AI visual assistant, and you are seeing a motion. Design a conversation between you and a person building a conversation about reasoning this motion. In conversations, you should indicate who said using ”User:”, ”AI:” in the beginning but these two words do not occur in sentences. The answers should be in a tone that an AI assistant is seeing the motion and answering the question. The scenario should always contain two people in the scene. Generate a conversation about building a story from two different motions. The flow of the conversation is as follows: 1. Creating a scenario. REMBER to make a story in this. 2. Reason about the motion or generate motion caption based on the scenario ”””Example: User: The current scene is [Two individuals sitting across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. Can you tell me what happened before? AI: Two people are doing arm-wrestling match, before that, two people will be doing fist dumps for fair play. User: Show me what will happen after that in motion format. AI: [One person is conducting a v-sign while the other stands still.]”””, ”””Example: User: Two friends are doing an arm-wrestling match, show me the motion of that. AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending left hand. They proceed to participate in a wrist-wrestling competition] User: Show me what happened before that in motion format. AI: [two people are doing fist dumps]. User: Why are they doing the fist dumps? AI: They are exchanging fist dumps to play a fair game in arm-wrestling.”””, ”””Example: User: The current scene is [Two individuals sitting across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. Can you tell me what happened before?. AI: Two people are doing arm-wrestling match, before that, two people will be doing fist dumps for fair play. User: The next scene is [One person is conducting a v-sign while the other stands still.]. Explain the reason for this motion. AI: After the arm-wrestling match, one person won the game. The person is showing this happiness to the audience.””” =========== Example format for the [motion caption]: - One person approaches, raises his/her right hand to grab the other person’s right forearm, places his/her left hand on it, and walks in the direction the grabbed person is facing. - Two people face each other, one person lifts his/her right leg and walks towards the other person, stopping half a meter away. - A person falls and braces himself/herself on the ground with his/her right hand. Another person approaches, squats down, and grabs his/her left arm with both hands to assist him/her in standing up. The content inside the bracket ([]) is a caption for the motion. This is for visualizing the motion, which is not given in textual form during inference. I will denote this as [motion caption]. Please denote [motion caption] when AI or the user has to answer in the motion sequence. Please make [motion caption] that is similar to the following action labels: [Action LABELS], and other motions like everyday routines (e.g., passing objects, greeting, communicating, etc.), and professional motions (e.g., Taekwondo, Latin dance, boxing, etc.) but still not necessary. Be creative too! Do not put [motion caption] in the same round, the user can also give motion to AI to reason from it too. Also, do not directly put [motion caption] twice in the round. You should put in only once, regarding both User and AI. [motion caption] are motion strings with skeleton information, which is for generating the motion. Do not repeat the caption. If you want to refer to these motions, just refer to it as the ’first motion’. But this motion string should be contained in the former to refer to. Try to make [motion caption] in details that do not require the previous context to generate the motion physically. ** Instead of the user fully describing what to do next, be more implicit, especially for the second motion, focusing more on the story. ** questions-answers not limited to the above examples. Questions should not be yes-no questions but wh-questions. The User-AI round should design at most 2. [motion caption] should appear only twice. Do not generate any new objects. Please follow the template from the example. It is better to keep the questions and answers concise. Try to be rational and keep in mind to make everything in sense, and the story smooth enough. Do not mention facial expressions or hands. Make the [motion caption] only ”twice” in the conversation. [motion caption] should always contain a description of two people. [motion caption] should have enough details for the motion, letting the model generate a correct motion by only accessing this caption without the previous context. Do not make the conversation more than three rounds.
@@
A.14 Prompts for LLM-Assisted Evaluation

### v5

@@ I Implementation details for Baselines
You are an AI visual assistant, and you are seeing a motion. Design a conversation between you and a person building a conversation about reasoning this motion. In conversations, you should indicate who said using ”User:”, ”AI:” in the beginning but these two words do not occur in sentences. The answers should be in a tone that an AI assistant is seeing the motion and answering the question. The scenario should always contain two people in the scene. Generate a conversation about building a story from two different motions. The flow of the conversation is as follows: 1. Creating a scenario. REMBER to make a story in this. 2. Reason about the motion or generate motion caption based on the scenario ”””Example: User: The current scene is [Two individuals sitting across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. Can you tell me what happened before? AI: Two people are doing arm-wrestling match, before that, two people will be doing fist dumps for fair play. User: Show me what will happen after that in motion format. AI: [One person is conducting a v-sign while the other stands still.]”””, ”””Example: User: Two friends are doing an arm-wrestling match, show me the motion of that. AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending left hand. They proceed to participate in a wrist-wrestling competition] User: Show me what happened before that in motion format. AI: [two people are doing fist dumps]. User: Why are they doing the fist dumps? AI: They are exchanging fist dumps to play a fair game in arm-wrestling.”””, ”””Example: User: The current scene is [Two individuals sitting across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. Can you tell me what happened before?. AI: Two people are doing arm-wrestling match, before that, two people will be doing fist dumps for fair play. User: The next scene is [One person is conducting a v-sign while the other stands still.]. Explain the reason for this motion. AI: After the arm-wrestling match, one person won the game. The person is showing this happiness to the audience.””” =========== Example format for the [motion caption]: - One person approaches, raises his/her right hand to grab the other person’s right forearm, places his/her left hand on it, and walks in the direction the grabbed person is facing. - Two people face each other, one person lifts his/her right leg and walks towards the other person, stopping half a meter away. - A person falls and braces himself/herself on the ground with his/her right hand. Another person approaches, squats down, and grabs his/her left arm with both hands to assist him/her in standing up. The content inside the bracket ([]) is a caption for the motion. This is for visualizing the motion, which is not given in textual form during inference. I will denote this as [motion caption]. Please denote [motion caption] when AI or the user has to answer in the motion sequence. Please make [motion caption] that is similar to the following action labels: [Action LABELS], and other motions like everyday routines (e.g., passing objects, greeting, communicating, etc.), and professional motions (e.g., Taekwondo, Latin dance, boxing, etc.)
but still not necessary. Be creative too! Do not put [motion caption] in the same round, the user can also give motion to AI to reason from it too. Also, do not directly put [motion caption] twice in the round. You should put in only once, regarding both User and AI. [motion caption] are motion strings with skeleton information, which are used to generate motion. Do not repeat the caption. If you want to refer to these motions, just refer to it as the ’first motion’. But this motion string should be contained in the former to refer to. Try to make [motion caption] in details that do not require the previous context to generate the motion physically. ** Instead of the user fully describing what to do next, be more implicit, especially for the second motion, focusing more on the story. ** questions-answers not limited to the above examples. Questions should not be yes-no questions but wh-questions. The User-AI round should design at most 2. [motion caption] should appear only twice. Do not generate any new objects. Please follow the template from the example. It is better to keep the questions and answers concise. Try to be rational and keep in mind to make everything in sense, and the story smooth enough. Do not mention facial expressions or hands. Make the [motion caption] only ”twice” in the conversation. [motion caption] should always contain a description of two people. [motion caption] should have enough details for the motion, letting the model generate a correct motion by only accessing this caption without the previous context. Do not make the conversation more than three rounds.
@@
S Prompts for LLM-Assisted Evaluation